Let ω be a k-form and let η be a 1-form. Find d(don η-ω^ dn)
Answer #1

d (dway) = dwn y + (1)kt how a dy = 0 - (-1) dwady d (wady) = dwa dy + (-uk wady • dwady to d (dan - wady) =-(-1) dwndy -